two person 27 kilometers apart start at the same time are together in 9 hours if they walk in the same direction but 3 hours if they walk in the opposite directions find out the rate of walking of each ?

Answer by ankor@dixie-net.com(22740)

two person 27 kilometers apart start at the same time are together in 9 hours if they walk in the same direction
but 3 hours if they walk in the opposite directions
find out the rate of walking of each ?
:
let a = the walking rate of the 1st person which is the fastest walker
let b = the walking rate of the 2nd person
:
Write a distance equation for each scenario
:
9(a - b) = 27; walking the same direction
3(a + b) = 27; walking in opposite directions
:
We can simplify both equations, divide the 1st eq by 9, the 2nd eq by 3:
a - b = 3
a + b = 9
-------------Adding eliminates b , find a
2a = 12
a = 12/2
a = 6 km/hr, 1st person's rate
then
6 + b = 9
b = 3 km/hr, 2nd person's rate
:
:
you can confirm these solutions in the both original equations